1/z(ez-1) at its poles


1
Expert's answer
2021-09-12T23:51:39-0400
"f(z)=\dfrac{1}{z(e^z-1)}"

The singularities are "z = 0" and "z = 2n\u03c0i, n = \u00b11, \u00b12, . . . ."


The singularity at "z = 0" is a pole of order 2 since "z = 0" is a zero of order 2 of "z(e^z-1)".


"\dfrac{1}{z(e^z-1)}=\dfrac{1}{z(z+z^2/2!+z^3/3!+...)}"

"=\dfrac{1}{z^2(1+z/2!+z^2/3!+...)}=\dfrac{1}{z^2}g(z), g(0)\not=0"

The singularities "z = 2n\u03c0i, n = \u00b11, \u00b12, . . . ." are simple poles since they are simple zeros of "\dfrac{1}{z(e^z-1)}."